# AI Agents vs. Agentic AI: A Conceptual Taxonomy, Applications and Challenges

### v3

@@ III The Emergence of Agentic AI from AI Agent Foundations
While AI Agents represent a significant leap in artificial intelligence capabilities, particularly in automating narrow tasks through tool-augmented reasoning, recent literature identifies notable limitations that constrain their scalability in complex, multi-step, or cooperative scenarios [137, 138, 139]. These constraints have catalyzed the development of a more advanced paradigm: Agentic AI. This emerging class of systems extends the capabilities of traditional agents by enabling multiple intelligent entities to collaboratively pursue goals through structured communication [140, 141, 142], shared memory [143, 144], and dynamic role assignment [14].
III-1 Conceptual Leap: From Isolated Tasks to Coordinated Systems
@@
Agentic AI systems represent an emergent class of intelligent architectures in which multiple specialized agents collaborate to achieve complex, high-level objectives [33]. As defined in recent frameworks, these systems are composed of modular agents each tasked with a distinct subcomponent of a broader goal and coordinated through either a centralized orchestrator or a decentralized protocol [16, 141]. This structure signifies a conceptual departure from the atomic, reactive behaviors typically observed in single-agent architectures, toward a form of system-level intelligence characterized by dynamic inter-agent collaboration.
A key enabler of this paradigm is goal decomposition, wherein a user-specified objective is automatically parsed and divided into smaller, manageable tasks by planning agents [39]. These subtasks are then distributed across the agent network. Multi-step reasoning and planning mechanisms facilitate the dynamic sequencing of these subtasks, allowing the system to adapt in real time to environmental shifts or partial task failures. This ensures robust task execution even under uncertainty [14].
Inter-agent communication is mediated through distributed communication channels, such as asynchronous messaging queues, shared memory buffers, or intermediate output exchanges, enabling coordination without necessitating continuous central oversight [147, 14]. Furthermore, reflective reasoning and memory systems allow agents to store context across multiple interactions, evaluate past decisions, and iteratively refine their strategies [148]. Collectively, these capabilities enable Agentic AI systems to exhibit flexible, adaptive, and collaborative intelligence that exceeds the operational limits of individual agents.
@@
In contrast, the right side of Figure 7 illustrates an Agentic AI system embedded in a comprehensive smart home ecosystem. Here, multiple specialized agents interact synergistically to manage diverse aspects such as weather forecasting, daily scheduling, energy pricing optimization, security monitoring, and backup power activation. These agents are not just reactive modules; they communicate dynamically, share memory states, and collaboratively align actions toward a high-level system goal (e.g., optimizing comfort, safety, and energy efficiency in real time). For instance, a weather forecast agent might signal upcoming heatwaves, prompting early pre-cooling via solar energy before peak pricing hours, as coordinated by an energy management agent. Simultaneously, the system might delay high-energy tasks or activate surveillance systems during occupant absence, integrating decisions across domains. This figure embodies the architectural and functional leap from task-specific automation to adaptive, orchestrated intelligence. The AI Agent acts as a deterministic component with limited scope, while Agentic AI reflects distributed intelligence, characterized by goal decomposition, inter-agent communication, and contextual adaptation, hallmarks of modern agentic AI frameworks.
@@
III-2 Key Differentiators between AI Agents and Agentic AI
To systematically capture the evolution from Generative AI to AI Agents and further to Agentic AI, we structure our comparative analysis around a foundational taxonomy where Generative AI serves as the baseline. While AI Agents and Agentic AI represent increasingly autonomous and interactive systems, both paradigms are fundamentally grounded in generative architectures, especially LLMs and LIMs. Consequently, each comparative table in this subsection includes Generative AI as a reference column to highlight how agentic behavior diverges and builds upon generative foundations.
A set of fundamental distinctions between AI Agents and Agentic AI particularly in terms of scope, autonomy, architectural composition, coordination strategy, and operational complexity are synthesized in Table I, derived from close analysis of prominent frameworks such as AutoGen [94] and ChatDev [149]. These comparisons provide a multi-dimensional view of how single-agent systems transition into coordinated, multi-agent ecosystems. Through the lens of generative capabilities, we trace the increasing sophistication in planning, communication, and adaptation that characterizes the shift toward Agentic AI.
TABLE I: Key Differences Between AI Agents and Agentic AI
@@ III-2 Key Differentiators between AI Agents and Agentic AI
Autonomy Level High autonomy within specific tasks. Higher autonomy with the ability to manage multi-step, complex tasks.
@@
Collaboration Operate independently. Involve multi-agent collaboration and information sharing.
@@
While Table I delineates the foundational and operational differences between AI Agents and Agentic AI, a more granular taxonomy is required to understand how these paradigms emerge from and relate to broader generative frameworks. Specifically, the conceptual and cognitive progression from static Generative AI systems to tool-augmented AI Agents, and further to collaborative Agentic AI ecosystems, necessitates an integrated comparative framework. This transition is not merely structural but also functional encompassing how initiation mechanisms, memory use, learning capacities, and orchestration strategies evolve across the agentic spectrum. Moreover, recent studies suggest the emergence of hybrid paradigms such as ”Generative Agents,” which blend generative modeling with modular task specialization, further complicating the agentic landscape. In order to capture these nuanced relationships, Table II synthesizes the key conceptual and cognitive dimensions across four archetypes: Generative AI, AI Agents, Agentic AI, and inferred Generative Agents. By positioning Generative AI as a baseline technology, this taxonomy highlights the scientific continuum that spans from passive content generation to interactive task execution and finally to autonomous, multi-agent orchestration. This multi-tiered lens is critical for understanding both the current capabilities and future trajectories of agentic intelligence across applied and theoretical domains.
TABLE II: Taxonomy Summary of AI Agent Paradigms: Conceptual and Cognitive Dimensions
@@
Memory Use No memory or short context window Optional memory or tool cache Shared episodic/task memory Subtask-local or contextual memory
@@
System Role Content generator Tool-using task executor Collaborative workflow orchestrator Subtask-level modular generator
To further operationalize the distinctions outlined in Table I, Tables III and II extend the comparative lens to encompass a broader spectrum of agent paradigms including AI Agents, Agentic AI, and emerging Generative Agents. Table III presents key architectural and behavioral attributes that highlight how each paradigm differs in terms of primary capabilities, planning scope, interaction style, learning dynamics, and evaluation criteria. AI Agents are optimized for discrete task execution with limited planning horizons and rely on supervised or rule-based learning mechanisms. In contrast, Agentic AI systems extend this capacity through multi-step planning, meta-learning, and inter-agent communication, positioning them for use in complex environments requiring autonomous goal setting and coordination. Generative Agents, as a more recent construct, inherit LLM-centric pretraining capabilities and excel in producing multimodal content creatively, yet they lack the proactive orchestration and state-persistent behaviors seen in Agentic AI systems.
The second table (Table III) provides a process-driven comparison across three agent categories: Generative AI, AI Agents, and Agentic AI. This framing emphasizes how functional pipelines evolve from prompt-driven single-model inference in Generative AI, to tool-augmented execution in AI Agents, and finally to orchestrated agent networks in Agentic AI. The structure column underscores this progression: from single LLMs to integrated toolchains and ultimately to distributed multi-agent systems. Access to external data, a key operational requirement for real-world utility, also increases in sophistication, from absent or optional in Generative AI to modular and coordinated in Agentic AI. Collectively, these comparative views reinforce that the evolution from generative to agentic paradigms is marked not just by increasing system complexity but also by deeper integration of autonomy, memory, and decision-making across multiple levels of abstraction.
TABLE III: Key Attributes of AI Agents, Agentic AI, and Generative Agents
@@
TABLE IV: Comparison of Generative AI, AI Agents, and Agentic AI
Feature Generative AI AI Agent Agentic AI
@@
Core Function Content generation Task-specific execution using tools Complex workflow automation
@@
Mechanism Prompt $\rightarrow$ LLM $\rightarrow$ Output Prompt $\rightarrow$ Tool Call $\rightarrow$ LLM $\rightarrow$ Output Goal $\rightarrow$ Agent Orchestration $\rightarrow$ Output
@@
Structure Single model LLM + tool(s) Multi-agent system
@@
External Data Access None (unless added) Via external APIs Coordinated multi-agent access
Key Trait Reactivity Tool-use Collaboration
@@
Furthermore, to provide a deeper multi-dimensional understanding of the evolving agentic landscape, Tables V through IX extend the comparative taxonomy to dissect five critical dimensions: core function and goal alignment, architectural composition, operational mechanism, scope and complexity, and interaction-autonomy dynamics. These dimensions serve to not only reinforce the structural differences between Generative AI, AI Agents, and Agentic AI, but also introduce an emergent category Generative Agents representing modular agents designed for embedded subtask-level generation within broader workflows [150]. Table V situates the three paradigms in terms of their overarching goals and functional intent. While Generative AI centers on prompt-driven content generation, AI Agents emphasize tool-based task execution, and Agentic AI systems orchestrate full-fledged workflows. This functional expansion is mirrored architecturally in Table VI, where the system design transitions from single-model reliance (in Generative AI) to multi-agent orchestration and shared memory utilization in Agentic AI. Table VII then outlines how these paradigms differ in their workflow execution pathways, highlighting the rise of inter-agent coordination and hierarchical communication as key drivers of agentic behavior.
@@
Furthermore, Table VIII explores the increasing scope and operational complexity handled by these systems ranging from isolated content generation to adaptive, multi-agent collaboration in dynamic environments. Finally, Table IX synthesizes the varying degrees of autonomy, interaction style, and decision-making granularity across the paradigms. These tables collectively establish a rigorous framework to classify and analyze agent-based AI systems, laying the groundwork for principled evaluation and future design of autonomous, intelligent, and collaborative agents operating at scale.
Each of the comparative tables presented from Table V through Table IX offers a layered analytical lens to isolate the distinguishing attributes of Generative AI, AI Agents, and Agentic AI, thereby grounding the conceptual taxonomy in concrete operational and architectural features. Table V, for instance, addresses the most fundamental layer of differentiation: core function and system goal. While Generative AI is narrowly focused on reactive content production conditioned on user prompts, AI Agents are characterized by their ability to perform targeted tasks using external tools. Agentic AI, by contrast, is defined by its ability to pursue high-level goals through the orchestration of multiple subagents each addressing a component of a broader workflow. This shift from output generation to workflow execution marks a critical inflection point in the evolution of autonomous systems.
@@
In Table VI, the architectural distinctions are made explicit, especially in terms of system composition and control logic. Generative AI relies on a single model with no built-in capability for tool use or delegation, whereas AI Agents combine language models with auxiliary APIs and interface mechanisms to augment functionality. Agentic AI extends this further by introducing multi-agent systems where collaboration, memory persistence, and orchestration protocols are central to the system’s operation. This expansion is crucial for enabling intelligent delegation, context preservation, and dynamic role assignment capabilities absent in both generative and single-agent systems. Likewise in Table VII dives deeper into how these systems function operationally, emphasizing differences in execution logic and information flow. Unlike Generative AI’s linear pipeline (prompt $\rightarrow$ output), AI Agents implement procedural mechanisms to incorporate tool responses mid-process. Agentic AI introduces recursive task reallocation and cross-agent messaging, thus facilitating emergent decision-making that cannot be captured by static LLM outputs alone. Table VIII further reinforces these distinctions by mapping each system’s capacity to handle task diversity, temporal scale, and operational robustness. Here, Agentic AI emerges as uniquely capable of supporting high-complexity goals that demand adaptive, multi-phase reasoning and execution strategies.
TABLE V: Comparison by Core Function and Goal
@@
TABLE VI: Comparison by Architectural Components
@@
TABLE VII: Comparison by Operational Mechanism
@@
TABLE VIII: Comparison by Scope and Complexity
@@
TABLE IX: Comparison by Interaction and Autonomy
@@
Furthermore, Table IX brings into sharp relief the operational and behavioral distinctions across Generative AI, AI Agents, and Agentic AI, with a particular focus on autonomy levels, interaction styles, and inter-agent coordination. Generative AI systems, typified by models such as GPT-3 [114] and and DALL·E https://openai.com/index/dall-e-3/, remain reactive generating content solely in response to prompts without maintaining persistent state or engaging in iterative reasoning. In contrast, AI Agents such as those constructed with LangChain[93] or MetaGPT[151], exhibit a higher degree of autonomy, capable of initiating external tool invocations and adapting behaviors within bounded tasks. However, their autonomy is typically confined to isolated task execution, lacking long-term state continuity or collaborative interaction.
Agentic AI systems mark a significant departure from these paradigms by introducing internal orchestration mechanisms and multi-agent collaboration frameworks. For example, platforms like AutoGen [94] and ChatDev [149] exemplify agentic coordination through task decomposition, role assignment, and recursive feedback loops. In AutoGen, one agent might serve as a planner while another retrieves information and a third synthesizes a report, each communicating through shared memory buffers and governed by an orchestrator agent that monitors dependencies and overall task progression. This structured coordination allows for more complex goal pursuit and flexible behavior in dynamic environments. Such architectures fundamentally shift the focus of intelligence from single-model outputs to emergent system-level behavior, wherein agents learn, negotiate, and update decisions based on evolving task states. Thus, the comparative taxonomy not only highlights increasing levels of operational independence but also illustrates how Agentic AI introduces novel paradigms of communication, memory integration, and decentralized control, paving the way for the next generation of autonomous systems with scalable, adaptive intelligence.

### v4

@@ III The Emergence of Agentic AI from AI Agent Foundations
While AI Agents represent a significant leap in artificial intelligence capabilities, particularly in automating narrow tasks through tool-augmented reasoning, recent literature identifies notable limitations that constrain their scalability in complex, dynamic, multi-step, and/or cooperative scenarios [143, 144, 145]. These constraints have catalyzed the development of a more advanced paradigm: Agentic AI. This emerging class of systems extends the capabilities of traditional AI Agents by enabling multiple intelligent entities to collaboratively pursue goals through structured communication [146, 147, 148], shared memory [149, 150], and dynamic role assignment [21].
III-1 Conceptual Leap: From Isolated Agents to Coordinated Systems
@@
Agentic AI systems represent an emergent class of intelligent architectures in which multiple specialized agents collaborate to achieve complex, high-level objectives utilizing collaborative reasoning and multi-step planning [40]. As defined in recent frameworks, these systems are composed of modular agents each tasked with a distinct subcomponent of a broader goal and coordinated through either a centralized orchestrator or a decentralized protocol [23, 147]. This structure signifies a conceptual departure from the individual, reactive behaviors typically observed in single-agent architectures, toward a form of system-level intelligence characterized by dynamic inter-agent collaboration.
A key enabler of this paradigm is goal decomposition, wherein a user-specified objective is automatically parsed and divided into smaller, manageable tasks by planning agents [46]. These subtasks are then distributed across the agent network. Multi-step reasoning and planning mechanisms facilitate the dynamic sequencing of these subtasks, allowing the system to adapt in real time to environmental changes or partial task failures. This agentic architecture ensures robust task execution even under uncertainty [21].
Inter-agent communication is mediated through distributed communication channels, such as asynchronous messaging queues, shared memory buffers, or intermediate output exchanges, enabling coordination without necessitating continuous central oversight [153, 21]. Furthermore, reflective reasoning and memory systems allow agents to store context across multiple interactions, evaluate past decisions, and iteratively refine their strategies [154]. Collectively, these capabilities enable Agentic AI systems to exhibit flexible, adaptive, cooperative, and collaborative intelligence that exceeds the operational limits of individual agents.
@@
In contrast, the right side of Figure 7 illustrates an Agentic AI system embedded in a comprehensive smart home ecosystem. Here, multiple specialized agents interact synergistically to manage diverse aspects such as weather forecasting, daily scheduling, energy pricing optimization, security monitoring, and backup power activation. These agents are not just reactive modules; they communicate dynamically, share memory states, and collaboratively align actions toward a high-level system goal (e.g., optimizing comfort, safety, and energy efficiency in real-time). For instance, a weather forecast agent might signal upcoming heatwaves, prompting early pre-cooling via solar energy before peak pricing hours, as coordinated by an energy management agent. Simultaneously, the system might delay high-energy tasks or activate surveillance systems during occupant absence, integrating decisions across domains. This figure embodies the architectural and functional leap from task-specific automation to adaptive, orchestrated intelligence. The AI Agent acts as a deterministic component with limited scope, while Agentic AI reflects distributed intelligence, characterized by goal decomposition, inter-agent communication, and contextual adaptation, demonstrating key characteristics of the modern agentic AI frameworks.
@@
III-2 Key Differences between AI Agents and Agentic AI
To systematically capture the evolution from Generative AI to AI Agents and further to Agentic AI, we structure our comparative analysis around a foundational taxonomy where Generative AI serves as the baseline. While AI Agents and Agentic AI systems represent increasingly autonomous and interactive systems, both paradigms utilize generative architectures as their foundations, especially LLMs and LIMs. Consequently, each comparative table in this subsection includes Generative AI as a reference column to highlight how agentic behavior builds on and then diverges from generative AI foundations.
A set of basic distinctions between AI Agents and Agentic AI systems, particularly in terms of scope, autonomy, architectural composition, coordination strategy, and operational complexity, are synthesized in Table I, which was derived from close analysis of prominent frameworks such as AutoGen [100] and ChatDev [155]. This comparison provides a multi-dimensional view of how single-agent systems transition into coordinated, multi-agent ecosystems. Through the perspective of generative capabilities, we trace the increasing sophistication in planning, communication, and adaptation that characterizes the shift toward Agentic AI systems.
TABLE I: Key Structural, Functional, and Operational Differences Between AI Agents and Agentic AI Systems. This table highlights the major distinctions between traditional AI Agents and more complex Agentic AI systems. It compares their definitions, levels of autonomy, capacity for handling task complexity, collaboration styles, learning and adaptation scope, and typical application domains. The comparison illustrates the evolution from task-specific, independently operating agents to coordinated, multi-agent systems capable of managing dynamic and large-scale workflows.
@@ III-2 Key Differences between AI Agents and Agentic AI
Autonomy Level High autonomy within specific tasks. Broad level of autonomy with the ability to manage multi-step, complex tasks and systems.
@@
Collaboration Operate independently. Involve multi-agent information sharing, collaboration and cooperation.
@@
While Table I delineates the foundational and operational differences between AI Agents and Agentic AI systems, a more granular taxonomy is required to understand how these paradigms emerge from and relate to broader generative AI frameworks. Specifically, the conceptual and cognitive progression from static Generative AI systems to tool-augmented AI Agents, and further to collaborative Agentic AI ecosystems, necessitates an integrated comparative framework. This transition is not merely structural but also functional encompassing how initiation mechanisms, memory use, learning capacities, and orchestration strategies evolve across the agentic spectrum. Moreover, recent studies suggest the emergence of hybrid paradigms such as ”Generative Agents,” which blend generative modeling with modular task specialization, further complicating the agentic AI landscape. In order to capture these nuanced relationships, Table II synthesizes the key conceptual and cognitive dimensions across four archetypes: Generative AI, AI Agents, Agentic AI systems, and inferred Generative Agents. By positioning Generative AI as a baseline technology, this taxonomy highlights the scientific, structural and application continuum that spans from passive content generation to interactive task execution and finally to autonomous, multi-agent orchestration. This multi-tiered perspective is critical for understanding both the current capabilities and future trends of agentic intelligence across theoretical and applied domains.
TABLE II: Summary of the Conceptual and Cognitive Taxonomy of AI Agent Paradigms Across Initiation, Adaptation, and Coordination Dimensions. This table synthesizes core cognitive and operational characteristics of four AI system types, comparing how they initiate tasks, adapt to goals, maintain temporal continuity, leverage memory, and coordinate actions. It captures the spectrum from stateless, prompt-driven Generative AI to highly coordinated and adaptive Agentic AI, while situating Generative Agents as modular components with localized generative capabilities within larger systems.
@@
Memory Use No memory or short context window Optional memory or tool cache Shared episodic/task memory Subtask-level or contextual memory
@@
Key Role Content generator Tool-based task executor Collaborative workflow orchestrator Subtask-level modular generator
To further operationalize the distinctions outlined in Table I, Tables III and II extend the comparison between agent paradigms to encompass a broader spectrum of paradigms including AI agents and agentic AI. Table III presents key architectural and behavioral attributes that highlight how each paradigm differs in terms of primary capabilities, planning scope, interaction style, learning dynamics, and evaluation criteria. As can be seen from the tables, AI Agents are optimized for discrete task execution with limited planning horizons and rely on supervised or rule-based learning mechanisms. In contrast, Agentic AI systems extend this capacity through multi-step planning, meta-learning, and inter-agent communication, positioning them for use in complex environments requiring autonomous goal setting and coordination. Generative Agents, as a more recent construct, inherit LLM-centric pretraining capabilities and excel in producing multi-modal content creation, yet they lack the proactive orchestration and state-persistent behaviors seen in Agentic AI systems.
The second table (Table III) provides a process-driven comparison across three agent categories: Generative AI, AI Agents, and Agentic AI. This framing emphasizes how functional pipelines evolve from prompt-driven single-model inference in Generative AI, to tool-augmented execution in AI Agents, and finally to orchestrated agent networks in Agentic AI systems. The structure column highlights this progression: from single LLMs to integrated tool-chains and ultimately to distributed multi-agent systems. Access to external data, a key operational requirement for real-world utility, also increases in sophistication, from absent or optional in Generative AI to modular and coordinated in Agentic AI. Collectively, these comparative views reinforce that the evolution from generative to agentic paradigms is marked not just by increasing system complexity but also by deeper integration of autonomy, memory, and decision-making across multiple levels of abstraction.
TABLE III: Key Differentiating Attributes of AI Agents, Agentic AI, and Generative Agents Across Capability, Learning, and Interaction Dimensions. This table outlines critical distinctions among three AI paradigms by comparing their primary capabilities, planning horizons, learning mechanisms, interaction styles, and evaluation criteria. It highlights the progression from task-specific execution in traditional AI Agents to the autonomous and adaptive behaviors of Agentic AI, and the creative, content-focused nature of Generative Agents.
@@
TABLE IV: Comparison of Generative AI, AI Agents, Agentic AI and Inferred Generative Agents Based on Core Function and Primary Goal. This table highlights the foundational purpose and operational focus of each system type, distinguishing their roles in AI workflows. It contrasts their core functions such as content generation, task execution, or workflow orchestration and clarifies the primary goals each category is optimized to achieve, from generating media to autonomously managing complex tasks.
Feature Generative AI AI Agent Generative Agent Agentic AI
@@
Core Function Content generation Task-specific execution using tools Simulated human-like behavior Complex workflow automation
@@
Mechanism Prompt $\rightarrow$ LLM $\rightarrow$ Output Prompt $\rightarrow$ Tool Call $\rightarrow$ LLM $\rightarrow$ Output Prompt $\rightarrow$ LLM + Memory/Planning $\rightarrow$ Output Goal $\rightarrow$ Agent Orchestration $\rightarrow$ Output
@@
Structure Single model LLM + tool(s) LLM + memory + behavior model Multi-agent system
@@
External Data Access None (unless added) Via external APIs Typically closed-world (simulated inputs) Coordinated multi-agent access
Key Trait Reactivity Tool-use Believability / Autonomy Collaboration
@@
Furthermore, to provide a deeper multi-dimensional understanding of the evolving agentic landscape, Tables V through IX extend the comparison over five critical dimensions: core function and goal alignment, architectural composition, operational mechanism, scope and complexity, and interaction-autonomy dynamics. These dimensions serve to not only reinforce the structural differences between Generative AI, AI Agents, and Agentic AI, but also introduce an emergent category of Generative Agents representing modular agents designed for embedded subtask-level generation within broader workflows [156]. Generative Agents are distinguished by their simulated human-like behavior, achieved through tightly integrated components such as language models, memory systems, and behavior planning modules, enabling them to operate believably and autonomously within typically closed-world environments. Table V situates the three paradigms in terms of their overarching goals and functional intent. While Generative AI centers on prompt-driven content generation, AI Agents emphasize tool-based task execution, and Agentic AI systems orchestrate full-fledged workflows. This functional expansion is mirrored architecturally in Table VI, where the system design transitions from single-model reliance (in Generative AI) to multi-agent orchestration and shared memory utilization in Agentic AI. Table VII then outlines how these paradigms differ in their workflow execution pathways, highlighting the rise of inter-agent coordination and hierarchical communication as key drivers of agentic behavior.
@@
Furthermore, Table VIII explores the increasing scope and operational complexity handled by these systems ranging from isolated content generation to adaptive, multi-agent collaboration in dynamic environments. Finally, Table IX synthesizes the varying degrees of autonomy, interaction style, and decision-making granularity across the paradigms. These tables collectively establish a rigorous framework to classify and analyze agent-based AI systems, laying the groundwork for theoretically grounded evaluation and future design of autonomous, intelligent, and collaborative agents operating at scale.
Tables V through IX offer a layered comparative analysis of Generative AI, AI Agents, and Agentic AI, anchoring the taxonomy in operational and architectural traits. Table V highlights core distinctions: Generative AI produces reactive content; AI Agents execute tool-based tasks; Agentic AI coordinates subagents for high-level workflow execution, marking a key shift in AI autonomy.
@@
In Table VI, the architectural distinctions are made explicit, especially in terms of system composition and control logic. Generative AI relies on a single model with no built-in capability for tool use or delegation, whereas AI Agents combine language models with auxiliary APIs and interface mechanisms to augment functionality. Agentic AI extends this further by introducing multi-agent systems where collaboration, memory persistence, and orchestration protocols are central to the system’s operation. This expansion is crucial for enabling intelligent delegation, context preservation, and dynamic role assignment capabilities absent in both generative and single-agent systems. Likewise in Table VII, differences in systems functionality and operation are presented, emphasizing distinctions in execution logic and information flow. Unlike Generative AI’s linear pipeline (prompt $\rightarrow$ output), AI Agents implement procedural mechanisms to incorporate tool responses mid-process. Agentic AI introduces recursive task reallocation and cross-agent messaging, thus facilitating emergent decision-making that cannot be captured by static LLM outputs alone. Table VIII further reinforces these distinctions by mapping each system’s capacity to handle task diversity, temporal scale, and operational robustness. Here, Agentic AI emerges as uniquely capable of supporting high-complexity goals that demand adaptive, multi-phase reasoning and execution strategies.
TABLE V: Comparison Generative AI, AI Agents, Agentic AI, and Inferred Generative Agents Based on Core Function and Primary Goal. This table highlights the foundational purpose and operational focus of each system type, distinguishing their roles in AI workflows. It contrasts their core functions such as content generation, task execution, or workflow orchestration and clarifies the primary goals each category is optimized to achieve, from generating media to autonomously managing complex tasks.
@@
TABLE VI: Comparison of Architectural Components Across Generative AI, AI Agents, Agentic AI, and Generative Agents. This table highlights key structural elements that define each AI paradigm, including core processing engines, prompt usage, tool and API integration, presence of multiple agents, and orchestration mechanisms. It illustrates the progression from single-model generative systems to complex multi-agent orchestration, while situating Generative Agents as modular units within broader workflows.
@@
TABLE VII: Comparison of Operational Mechanisms Among Generative AI, AI Agents, Agentic AI, and Generative Agents. This table details the driving forces behind each system’s operation, their modes of interaction, approaches to workflow management, and patterns of information flow. It highlights the transition from reactive, single-step generation in Generative AI to coordinated multi-agent workflows in Agentic AI, with Generative Agents functioning as modular contributors within broader task sequences.
@@
TABLE VIII: Comparison of Task Scope and Complexity Across Generative AI, AI Agents, Agentic AI, and Generative Agents. This table examines the breadth and difficulty of tasks each AI paradigm typically handles, illustrating the shift from generating single content pieces to managing complex, multi-agent workflows.
@@
TABLE IX: Comparison of Interaction and Autonomy Levels Among Generative AI, AI Agents, Agentic AI, and Generative Agents. This table analyzes varying degrees of autonomy, modes of external and internal interaction, and decision-making processes across AI paradigms. It highlights the progression from prompt-dependent Generative AI to highly autonomous, multi-agent coordination in Agentic AI, with Generative Agents operating autonomously within limited sub-task scopes.
@@
Furthermore, Table IX highlights the operational and behavioral distinctions across Generative AI, AI Agents, and Agentic AI, with a particular focus on autonomy levels, interaction styles, and inter-agent coordination. Generative AI models such as GPT-3 [120] and DALL·E https://openai.com/index/dall-e-3/, remain reactive generating content solely in response to prompts without maintaining a persistent state or engaging in iterative reasoning. In contrast, AI Agents such as those constructed with LangChain[99] or MetaGPT[157], exhibit a higher degree of autonomy, capable of initiating external tool invocations and adapting behaviors within bounded tasks. However, their autonomy is typically confined to isolated task execution, lacking long-term state continuity or collaborative interaction.
Agentic AI systems mark a significant departure from these paradigms by introducing internal orchestration mechanisms and multi-agent collaboration frameworks. For example, platforms like AutoGen [100] and ChatDev [155] exemplify agentic coordination through task decomposition, role assignment, and recursive feedback loops. In AutoGen, one agent might serve as a planner while another retrieves information and a third synthesizes a report, each communicating through shared memory buffers and governed by an orchestrator agent that monitors dependencies and overall task progression. This structured coordination allows for more complex goal pursuit and flexible behavior in dynamic environments. Such architectures fundamentally shift the focus of intelligence from single-model-based outputs to system-level behavior, wherein agents learn, adapt, and update decisions based on evolving task states. Thus, this comparative taxonomy not only highlights increasing levels of operational independence but also illustrates how Agentic AI introduces novel paradigms of communication, memory integration, and decentralized control, paving the way for the next generation of autonomous systems with scalable, adaptive intelligence.